An object has an initial velocity of 20 m/s how long is it accelerate at a constant rate of 4 m/s squared before the speed equals three times the initial speed


1
Expert's answer
2021-08-16T08:34:06-0400

Given:

"v_i=20\:\rm m/s"

"v_f=60\:\rm m/s"

"a=4\:\rm m/s^2"


The time of accelerating

"t=\frac{v_f-v_i}{a}=\frac{60\:\rm m/s-20\:\rm m/s}{4\:\rm m/s^2}=10\:\rm s"